# A New Approach to the Representation Theory of Lorentzian Pseudo-Tensors

### v1

@@ II.4 The Effect of Parity and Time Reversal on Boosts and Rotations
All this is merely stating that we know from the velocity dependence of boosts, under a spacial or temporal reflection, the boost parameter will pick up a sign. In the representation theoretic language introduced above, we would say boost parameters (rapidities, velocities etc.) are charged under the $PT$ representation of the reflection group. [^6] This point of view is further defended in Sec. (II.6.1). With an understanding of how $\mathcal{P}$ and $\mathcal{T}$ interact with boosts, one may be curious about rotations. To make some vector co-linear with a specified axis, suppose we act with some rotation matrix. After this we may apply a time reversal or parity reflection. It is straightforward to convince oneself that were we to reflect in either time or space first, the rotation matrix required to end up in the same state resulting from the original order of operations, is simply the rotation matrix we started with. [^7] If the vector is odd under $\mathcal{T}$, only the application the same rotation matrix will make it so that time-reflection followed by rotation agrees with rotation followed by time-reflection. If the vector is odd under $\mathcal{P}$, any plane of rotation will have both axes flipped, and so can be understood to have merely been rotated through $\pi$, and $2D$ rotations in the plane commute. If the vector is even under either reflection, the commutation is trivial. This can also be understood via commutation relations:
@@
We might also say that the angle parameter $\theta$ is uncharged under the discrete reflection group, i.e. $K[\theta]=1$. It is also well known that any proper orthochronus Lorentz transform may be decomposed as a pure boost, followed by a rotation:
@@
so we may condense the commutation relations:
@@
As discussed in Appendix (B), the reflection of all boost parameters is equivalent to saying we have acted via an outer automorphism $\mathcal{\hat{O}}$ upon our Lie group:
This is precisely the behaviour expected from a semi-direct product (see Appendix (A)): commuting a first kind of group element past a second kind, results in an automorphism being applied to the first kind of group element. This observation tells us explicitly how to build the semi-direct product. We are seeking a map $\psi$ that takes elements of the discrete reflection group, and maps them to automorphisms of $\operatorname{SO}^{+}(1,3)$:
Let $\psi$ be the following map:
II.5 Constraining Reflection Charges with Lorentz Transforms
With the reasonable demand that all observers agree about the representations objects take under $\operatorname{O}(1,3)$, we find some much needed restrictions on the kinds of objects which can enter our theories. Given a four vector $V^{\mu}$ in some frame, with each Cartesian component having arbitrary charges $A,B,C,D$ under the discrete reflection group:
$K[V^{\mu}]=\begin{pmatrix}A\\ B\\ C\\ D\end{pmatrix}$ (22)
then another frame rotated from the first, by an angle $\theta$ in the $xy$ plane, will see the vector with charges
@@ II.5 Constraining Reflection Charges with Lorentz Transforms
$K[V^{\mu}]=\begin{pmatrix}A\\ \mathbf{B}\end{pmatrix}.$ (24)
@@
As previously discussed, the rapidity $\omega\propto v$ is charged under $PT$. As such, $K[\cosh\omega]=1$ and $K[\sinh\omega]=PT$ which tells us the charges in the new frame are
@@
For all observers to agree about the image of the vector under the discrete reflection group, we must have that $B\cdot PT=A$. This give us only four kinds of four-vector which have well defined physical transformation properties for all observers related by $\operatorname{O}(1,3)$. These are:
@@
We will name these respectively ‘coordinate type’ (c), ‘momentum type’ (m), ‘axial type’ (a), and ‘polarization type’ (p) four vectors. When acting on these particular vector representations we will denote the operators with the shorthand $\mathcal{P}_{(c)}$, $\mathcal{T}_{(m)}$ etc, and utilize the notation that $K[x^{\mu}]=(c)$, $K[p^{\mu}]=(m)$, etc. It is convenient to note that the basic operators here are all related up to a sign:
@@
II.6 Simplest Tensor Representations of $\operatorname{O}(1,3)$
II.6.1 Generators of Lorentz Transforms
@@
This is necessarily the representation which the metric tensor lives in. Furthermore, it is well known that the adjoint representation of the (proper orthochronus) Lorentz group lives in the anti-symmetric part of this tensor product space. Thus as the generators of boosts live in the $0i$ components of this kind of tensor, and the generators of rotations live in the $ij$ components, this serves to show that the generators of boosts and rotations are charged under discrete reflection group as $PT$ and $1$, respectively. Let us confirm that the operations of $\mathcal{P}$ and $\mathcal{T}$ in the defining representation, acting on this tensor are the same as one another, as we anticipate from the charges. With
$g(\mathbf{K},\mathbf{L})=\begin{pmatrix}0&K_{x}&K_{y}&K_{z}\\ -K_{x}&0&L_{z}&-L_{y}\\ -K_{y}&-L_{z}&0&L_{x}\\ -K_{z}&L_{y}&-L_{x}&0\end{pmatrix},$ (30)
we can see
$\begin{split}&\mathcal{P}_{(c)}\>g(\mathbf{K},\mathbf{L})\mathcal{P}_{(c)}=\left(\mathcal{-P}_{(c)}\right)g(\mathbf{K},\mathbf{L})\left(\mathcal{-P}_{(c)}\right)\\ =&\vspace{1ex}\mathcal{T}_{(c)}\>g(\mathbf{K},\mathbf{L})\mathcal{T}_{(c)}\\ =&\vspace{1ex}\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}\begin{pmatrix}0&K_{x}&K_{y}&K_{z}\\ -K_{x}&0&L_{z}&-L_{y}\\ -K_{y}&-L_{z}&0&L_{x}\\ -K_{z}&L_{y}&-L_{x}&0\end{pmatrix}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\times\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}\\ =&\begin{pmatrix}0&-K_{x}&-K_{y}&-K_{z}\\ K_{x}&0&L_{z}&-L_{y}\\ K_{y}&-L_{z}&0&L_{x}\\ K_{z}&L_{y}&-L_{x}&0\end{pmatrix}=g(-\mathbf{K},\mathbf{L}).\end{split}$ (31)
This all matches what we have understood so far, and is as far as most textbooks might go to inspecting the action of these operators. The cautious among us may be concerned that the generators of angular momentum have the same transformation properties under time reversal and parity, and that the transformation is trivial. Of course this is not a problem so long as the angular momentum itself has the correct transformation properties.
II.6.2 Relativistic Angular Momentum
Let us consider the representation of the relativistic angular momentum tensor, given by:
@@ II.6.2 Relativistic Angular Momentum
Here $\mathbf{N}$ is the dynamic mass moment vector and $\mathbf{J}$ is the angular momentum vector. The representation of this tensor under the discrete reflection group is the tensor product of two vectors, one coordinate type and one momentum type:
@@
From the assignments of charges to the components we can see this is precisely what is expected: the angular momentum three vector will be odd under time reversal and even under parity. From the point of view of application of operators this case is the most interesting thus far, since we are now combining vectors under different representations of $\operatorname{O}(1,3)$. Firstly, note the representation of $\mathcal{P}$ on $p^{\mu}$ and on $x^{\mu}$ is the same: $\mathcal{P}_{(c)}=\mathcal{P}_{(m)}$, and so the parity operator will act just as it did with the boost and rotation generators:
@@
This result shows that we may think of the tensor as having two indices which transform differently under the discrete reflection group, analogous to how co- and contravariant indices of the same tensor transform differently. However it is clear in this case the transform is ambivalent about ‘which’ index transforms which way: we get the same result of the transform regardless. As such we may think of there being unambiguous tensor representations of the discrete reflection group, and we might say $K[M^{\alpha\beta}]=(cm)$, or (recall $g$ from above) $K[g]=(cc)$. [^8] These ‘vector rep’ indicators of tensor product reps are in fact not unique, since for example $(cc)$ and $(mm)$ are the same representation (as well as $(ca)$ and $(mp)$ etc.) A more satisfying representation theory of generic Lorentz tensors under $\operatorname{K}_{4}$ is presented in the Sec. (II.7). Regardless of the redundancy in this section, the important point is that these make the transformation properties of the tensor unambiguous. From this, we know precisely which representation of the parity and time reversal operators to act upon the tensor with, and we know in fact the order in which they act does not matter. Now we may correctly and unambiguously perform the time reversal of an angular momentum tensor.
$\begin{split}&\mathcal{T}\left[M(\mathbf{N},\mathbf{J})\right]\\ \vspace{0.8em}=&\mathcal{T}_{c}\>M(\mathbf{N},\mathbf{J})\mathcal{T}_{m}\\ \vspace{0.8em}=&\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}\begin{pmatrix}0&N_{x}&N_{y}&N_{z}\\ -N_{x}&0&J_{z}&-J_{y}\\ -N_{y}&-J_{z}&0&J_{x}\\ -N_{z}&J_{y}&-J_{x}&0\end{pmatrix}\\ &\qquad\qquad\qquad\qquad\qquad\times\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\\ \end{pmatrix}\\ =&\begin{pmatrix}0&N_{x}&N_{y}&N_{z}\\ -N_{x}&0&-J_{z}&J_{y}\\ -N_{y}&J_{z}&0&-J_{x}\\ -N_{z}&-J_{y}&J_{x}&0\end{pmatrix}=M(\mathbf{N},-\mathbf{J}).\end{split}$ (37)
@@
II.6.3 The Faraday Tensor and the Potential $A^{\mu}$
While $A^{\mu}$ is not a 2-tensor, a discussion of the Faraday tensor is incomplete without addressing it first. The parity inversion properties of $A^{\mu}$ are rarely up for debate. It is the time reversal of $A^{\mu}$ which is the cause of contention in the literature. In particular, it is the question of whether a time reversal should ‘include’ a charge conjugation. This debate is due in part to the Feynman–Stückelberg interpretation of anti-particle solutions, and in part due to the anti-unitarity of the time reversal operator. However, so long as we are considering the case of classical fields, we can cleanly separate time reversal from charge conjugation without debate. In the following the time reversal of a stationary charge $q$, will be understood to produce a description of an identical system, and not a description of a system of a stationary charge of $-q$. This can be stated as the postulate that classical electrical charges live in the trivial representation of the discrete reflection group.
It is not clear a priori what type of vector $A^{\mu}$ is. One can argue from its equations of motion that it should have the same properties as a current four vector. However this requires first assuming those equations are parity and time reversal invariant. The answer can be gleaned from the much less assumptive but important expectation that the canonical momentum of the electromagnetic field has well defined transformation properties under the discrete reflection group for all observers. The momentum which enters the Hamiltonian of the electromagnetic field is given by JACKSON3RD :
$P^{\mu}=\left(p^{\mu}+eA^{\mu}\right),$ (38)
where $p^{\mu}$ is the mechanical momentum of the relevant particle with charge $e$. If observers related by (proper orthochronus) Lorentz transforms are to agree about the charges of $P^{\mu}$ under the discrete reflection group, it must be the case that $K[eA^{\mu}]=K[p^{\mu}]=(m)$. As we are assuming $e$ is invariant under the reflection group, we conclude $K[A^{\mu}]=K[p^{\mu}]=1\oplus\vec{PT}=(m)$, i.e. $A^{\mu}$ is a momentum type vector.
This takes an identical representation to the angular momentum tensor:
$K[F^{\mu\nu}]=\begin{pmatrix}T\\ \vec{P}\end{pmatrix}\otimes\begin{pmatrix}1\\ \vec{PT}\end{pmatrix}=\begin{pmatrix}T&\vec{P}\\ \vec{P}&T\mathbb{I}_{3}\end{pmatrix},$ (40)
and so of course parity and time reversal will act the same here as they did upon the angular momentum tensor. I.e.:
This is what is taught in most physics textbooks, and finally we can see a consistent representation theoretic defense of it. The only room for disagreement as far as the author can surmise, is the debate of whether or not charges should take non-trivial representations of the discrete reflection group $\operatorname{K}_{4}$, which presently seems a discussion more appropriate to be had at the quantum mechanical level. We can also relate the computation of the time reversal to the parity inverted state, and see clearly why what Arntzenius & Greaves GREAVESARNTZENIUS dubbed ‘Malaments proposal’ [^9] That the time reversal of the Faraday tensor should look like the application of parity, followed by an overall additional minus sign on the whole tensor. works:
II.7 Generic Tensor Representations of $\operatorname{O}(1,3)$
We are now in a position to build up to a rather satisfying conclusion regarding generic tensor representations of $\operatorname{O}(1,3)$. We will see for every tensor power, the pattern of there being precisely four representations under $\operatorname{K}_{4}$ will persist: these can then be uniquely and consistently named so that the tensor product of generic representations simultaneously follows the representation theory of each $\operatorname{SO}^{+}(1,3)$ and $\operatorname{K}_{4}$.
II.7.1 Scalars
The case of scalars under $\operatorname{O}(1,3)$ is already accounted for by the one dimensional irreducible representations of $\operatorname{K}_{4}$. They are what we have called $1,P,T,PT$.
II.7.2 Vectors
There are precisely four well defined vector representations of $\operatorname{O}(1,3)$. Previously these were named $(c)$, $(m)$, $(a)$, and $(p)$. However one could imagine a more efficient naming scheme: let us instead label our vector representations by the sign difference $\mathcal{P}$ and $\mathcal{T}$ take from the defining (coordinate) representation, given in Tab. (2).

### v2

@@ II.3 The Effects of Parity and Time Reversal on Boosts and Rotations
That is to say under a spacial or temporal reflection the boost parameter will pick up a sign, just as a velocity will. In the language introduced above, we would say boost parameters (rapidities, velocities etc.) are charged under the $PT$ representation of the reflection group. With an understanding of how $\mathcal{P}$ and $\mathcal{T}$ interact with boosts, next we turn to rotations. It turns out rotations commute with parity and time reversal. As a generic example, suppose one hopes to make some vector co-linear with a specified target direction, via a rotation. In 3D any rotation can be understood as a rotation in a single plane, i.e. around a single axis; the plane in this case is the span of the original vector and the target direction. If the vector is odd under $\mathcal{T}$, the negation of the vector and the target vector is equivalent to a rotation by $\pi$ in that plane, and all rotations in the plane commute. If the vector is odd under $\mathcal{P}$, the plane of rotation will have both axes flipped, which once again may be understood as a rotation through $\pi$. If the vector is even under either reflection, the commutation is trivial. Written in the same form as the above, we have:
@@
We say the rotation parameter $\theta$ is uncharged under the discrete reflection group: $K[\theta]=1$. It is well known that any proper orthochronus Lorentz transform may be decomposed as a pure boost, followed by a rotation:
@@
so we may condense the commutation relations: [^6] Though not needed on a first read, appendix. (B.3) continues directly from the current subsection.
@@
II.4 Constraining Reflection Charges with Lorentz Transforms
The demand that vectors (and tensorial objects generally) represent covariant physical quantities is tantamount to the declaration that all inertial observers agree about the representations objects take under $\operatorname{O}(1,3)$, including reflections. Without this demand, naively each component of a vector could be assigned a distinct reflection charge leading to $4^{4}$ possible vector representations, most of which would be constructions which have preferred frames. As such the demand of covariance provides a much needed restriction on the kinds of objects which can enter our theories. Take a four vector $V^{\mu}$ in some inertial frame, with each Cartesian component having arbitrary charges $A,B,C,D$ under the discrete reflection group
$K[V^{\mu}]=\begin{pmatrix}A\\ B\\ C\\ D\end{pmatrix}.$ (15)
Another frame rotated from the first by an angle $\theta$ in the $xy$ plane will see the vector with charges
@@ II.4 Constraining Reflection Charges with Lorentz Transforms
$K[V^{\mu}]=\begin{pmatrix}A\\ \vec{B}\end{pmatrix},$ (17)
@@
As previously discussed, the rapidity $\omega\propto v$ has reflection charge $PT$. As such, $K[\cosh\omega]=1$ and $K[\sinh\omega]=PT$ which tells us the charges in the new frame are
@@
For this transform to be covariant under the reflections, we must have that $B\cdot PT=A$. This give us only four kinds of four-vector which have well defined physical transformation properties for all observers related by $\operatorname{O}(1,3)$. These are:
@@
We will name these respectively ‘coordinate type’ (c), ‘momentum type’ (m), ‘axial type’ (a), and ‘polarization type’ (p) four-vectors. When acting on these particular vector representations we will denote the operators with the shorthand $\mathcal{P}_{(c)}$, $\mathcal{T}_{(m)}$, etc, and utilize the notation $K[x^{\mu}]=(c)$, $K[p^{\mu}]=(m)$, etc. It is convenient to note that the basic operators here are all related up to a sign:
@@
With vectors classified, we can begin to ask about the parity and time reversal of tensors generally.
II.5 Simple 2-Tensor Representations of $\operatorname{O}(1,3)$
II.5.1 Generators of Lorentz Transforms
@@
This is the representation which the metric tensor lives in, since the metric components are defined as the inner product of coordinate basis vectors. From the information given in Sec. (II.3) and standard representation theory, it can also be derived that the generators of Lorentz transforms upon vectors, live in this representation. The generators of boosts live in the $0i$ components of this kind of tensor, and the generators of rotations live in the $ij$ components. This agrees with our understanding that the generators of boosts and rotations are charged under discrete reflection group as $PT$ and $1$, respectively.
Define shorthand for a generic generator with boost parameters $\mathbf{K}$ and rotation parameters $\mathbf{L}$
$g(\mathbf{K},\mathbf{L})=\begin{pmatrix}0&K_{x}&K_{y}&K_{z}\\ K_{x}&0&L_{z}&-L_{y}\\ K_{y}&-L_{z}&0&L_{x}\\ K_{z}&L_{y}&-L_{x}&0\end{pmatrix}.$ (23)
This tensor was constructed out of coordinate type vectors, so let us confirm that the operations of $\mathcal{P}$ and $\mathcal{T}$ acting in the defining representation behave as we anticipate. For time reversal we have
$\begin{split}&\mathcal{T}\left[g(\mathbf{K},\mathbf{L})\right]=\vspace{1ex}\mathcal{T}_{(c)}\>g(\mathbf{K},\mathbf{L})\mathcal{T}_{(c)}\\ =&\vspace{1ex}\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}\begin{pmatrix}0&K_{x}&K_{y}&K_{z}\\ K_{x}&0&L_{z}&-L_{y}\\ K_{y}&-L_{z}&0&L_{x}\\ K_{z}&L_{y}&-L_{x}&0\end{pmatrix}\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}\end{split}$
$\>=\begin{pmatrix}0&-K_{x}&-K_{y}&-K_{z}\\ -K_{x}&0&L_{z}&-L_{y}\\ -K_{y}&-L_{z}&0&L_{x}\\ -K_{z}&L_{y}&-L_{x}&0\end{pmatrix}=g(-\mathbf{K},\mathbf{L}).$ (24)
As for parity we have that $\mathcal{P}_{(c)}=-\mathcal{T}_{(c)}$, and so the signs can be made to cancel and the calculation gives the same: $\mathcal{P}\left[g(\mathbf{K},\mathbf{L})\right]=g(-\mathbf{K},\mathbf{L})$. We see both reflections are in agreement with the charges assigned in Eq. (22). The cautious among us may be concerned that the generators of angular momentum have the same transformation properties under time reversal and parity, and that the transformation is trivial. Of course this is not a problem so long as the angular momentum itself has the correct transformation properties.
II.5.2 Relativistic Angular Momentum
Consider a relativistic angular momentum tensor $M$:
@@ II.5.2 Relativistic Angular Momentum
Here $\mathbf{N}$ is the dynamic mass moment vector and $\mathbf{J}$ is the angular momentum vector. This tensor looks superficially identical to the previous example. However this tensor is built from vectors of different representations under the discrete reflection group, one coordinate type and one momentum type:
@@
From the assignments of charges to the components we can see this is precisely what is expected: the angular momentum three vector will be odd under time reversal and even under parity. This case is the most interesting thus far, since we are combining vectors under different representations of $\operatorname{O}(1,3)$. Note the representation of $\mathcal{P}$ on $p^{\mu}$ and on $x^{\mu}$ is the same: $\mathcal{P}_{(c)}=\mathcal{P}_{(m)}$, and so the parity operator will act just as it did with the boost and rotation generators:
@@
This result shows that we may think of the tensor as having two indices which transform differently under the discrete reflection group, analogous to how co- and contravariant indices of the same tensor transform differently. However it is clear in this case the transform is ambivalent about ‘which’ index transforms which way: we get the same result of the transform regardless. As such we may think of there being unambiguous tensor representations of the discrete reflection group, and we might say $K[M^{\alpha\beta}]=(cm)$, or (recall $g$ from above) $K[g]=(cc)$. [^7] These ‘vector rep’ indicators of tensor product reps are in fact not unique, since for example $(cc)$ and $(mm)$ are the same representation (as well as $(ca)$ and $(mp)$ etc.) A more satisfying representation theory of generic Lorentz tensors under $\operatorname{K}_{4}$ is presented in the Sec. (II.6). Regardless of the redundancy in this section, the important point is that these make the transformation properties of the tensor unambiguous. From this, we know precisely which representation of the parity and time reversal operators to act upon the tensor with, and we know in fact the order in which they act does not matter. Now we may correctly and unambiguously perform the time reversal of an angular momentum tensor.
$\begin{split}&\mathcal{T}\left[M(\mathbf{N},\mathbf{J})\right]=\mathcal{T}_{c}\>M(\mathbf{N},\mathbf{J})\mathcal{T}_{m}\\ \vspace{0.8em}=&\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}\begin{pmatrix}0&N_{x}&N_{y}&N_{z}\\ -N_{x}&0&J_{z}&-J_{y}\\ -N_{y}&-J_{z}&0&J_{x}\\ -N_{z}&J_{y}&-J_{x}&0\end{pmatrix}\\ &\qquad\qquad\qquad\qquad\qquad\times\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\\ \end{pmatrix}\\ =&\begin{pmatrix}0&N_{x}&N_{y}&N_{z}\\ -N_{x}&0&-J_{z}&J_{y}\\ -N_{y}&J_{z}&0&-J_{x}\\ -N_{z}&-J_{y}&J_{x}&0\end{pmatrix}=M(\mathbf{N},-\mathbf{J}).\end{split}$ (30)
@@
II.6 Generic Tensor Representations of $\operatorname{O}(1,3)$
We are now in a position to build up to a rather satisfying conclusion regarding generic tensor representations of $\operatorname{O}(1,3)$. We will see for every tensor power of the fundamental vector represenation of $\operatorname{SO}^{+}(1,3)$, the pattern of there being precisely four representations under $\operatorname{K}_{4}$ will persist: these can then be uniquely and consistently named so that the tensor product of generic representations simultaneously follows the representation theory of each $\operatorname{SO}^{+}(1,3)$ and $\operatorname{K}_{4}$.
II.6.1 Scalars and Vectors
The case of scalars under $\operatorname{O}(1,3)$ is already accounted for by the one dimensional irreducible representations of $\operatorname{K}_{4}$. They are what we have called $1,P,T,PT$. We have also seen there are precisely four well defined vector representations of $\operatorname{O}(1,3)$. Previously these were named $(c)$, $(m)$, $(a)$, and $(p)$. However one could imagine a more direct naming scheme: let us instead label our vector representations by the sign difference $\mathcal{P}$ and $\mathcal{T}$ take from the defining (coordinate) representation, given in Tab. (2).